# What is the nature of late infall streamers?

### v1

@@
What is the nature of late infall streamers?
@@
(October 28, 2025)
Growing observational evidence suggests that Class II protoplanetary disks may undergo substantial interactions with their environment in the form of late infall. This mass inflow predominantly manifests itself in the form of so-called streamers: filaments and arcs of gas connecting large-scale, extended gas structures to disk scales. Prevalent late infall has far-reaching consequences for planet formation theory, challenging the long-standing treatment of evolved disks in isolation. In this work, we investigate the emergence of late-infall streamers in different formation scenarios, their morphology and multiplicity, as well as their dependence on environmental conditions. We conduct this investigation by performing 3D hydrodynamical simulation using the grid-based code FARGO3D, which we post-process to obtain synthetic observations using the Monte Carlo radiative transfer code RADMC3D. We find that, while a late infall event in the form of a single encounter with a ”cloudlet” of gas can produce a streamer via an interplay between the fall-back of bound material and shocks, such features dissipate quickly, on a timescale of ${\sim}$10\text{\,}\mathrm{kyr}$\$. Furthermore, we find that streamers can also form naturally in a turbulent, dense environment without the need for such encounters, which could act to reconcile short-lived streamers with ubiquitous detection of these structures. Here, we find multiple co-existing streamers for a disk velocity relative to the interstellar medium of $v_{\mathrm{sys}}=$0.5\text{\,}\mathrm{km}\text{\,}{\mathrm{s}}^{-1}$\$ and a turbulent velocity dispersion of $\sigma_{\mathrm{turb}}=$0.5\text{\,}\mathrm{km}\text{\,}{\mathrm{s}}^{-1}$\$, in which case an angular momentum flux of $\dot{J}\sim${10}^{38}\text{\,}\mathrm{g}\text{\,}{\mathrm{cm}}^{2}\text{\,}{\mathrm{s}}^{-2}$\$ arises. We find considerable dependence of the streamer morphology on the environment, which may act as a utility to constrain the physical conditions of the gas surrounding planet-forming disk, and therefore the conditions under which planets form.
@@ 1 Introduction
Planet formation theory has evolved greatly in recent years, with more complete models now connecting the large range of size scales of dust grain growth, the formation of planetary cores, migration and gas accretion (for a review, see Drążkowska et al. 2023). Despite the steady advancement, protoplanetary disks, the birthplaces of planets, are still largely treated as being isolated from their environment. After their initial formation, the dust budget is treated as fixed, being converted into planetary cores through internal processes. Issues remain in this framework; in particular, recent observations indicate that the dust mass budget of fully formed protoplanetary disks is insufficient to form the giant planets we observe (Manara et al. 2018).
@@
Understanding the mechanisms that can create streamer-like structures and shape their morphology, as well as identifying their potential origin, is vital for our understanding of the dynamical and mass-budget evolution of protoplanetary disks, with potential far-reaching consequences for planet formation as a whole. In this work, we investigate two different types of interactions of fully formed protoplanetary disks with their environment using hydrodynamical simulations. First, we probe how a single, well-characterized interaction with a cloudlet of gas can create various streamer-like observational signatures (Section 3.1). We then consider a disk that is embedded in the turbulent interstellar medium to examine structures that arise in the absence of a distinct source of infall (Section 3.2). While signatures of single interactions are not long-lived, those arising from constant interactions with the ISM could persist over longer time scales and would therefore be much more abundantly observable.
@@ 2 Methods
We performed 3D hydrodynamical simulations on a spherical grid with logarithmic spacing in the radial direction, and uniform spacing in the polar and azimuthal direction. The computations were performed using the FARGO3D code (Benítez-Llambay & Masset 2016), capable of GPU multiprocessing while utilizing the advection algorithm by Masset (2000). The radial extent of the grid is from $r_{\mathrm{min}}=$10\text{\,}\mathrm{AU}$\$ to $r_{\mathrm{max}}=$5000\text{\,}\mathrm{AU}$\$, except the simulation presented in Section 3.1, where $r_{\mathrm{min}}=$5\text{\,}\mathrm{AU}$\$. We moved the outer edge further outward to $r_{\mathrm{max}}=$50\,000\text{\,}\mathrm{AU}$\$ for simulations with a turbulent initial condition, described in later sections. In the polar direction, the range is from $\theta_{\mathrm{min}}=$10\text{\,}\mathrm{\SIUnitSymbolDegree}$\$ to $\theta_{\mathrm{max}}=$170\text{\,}\mathrm{\SIUnitSymbolDegree}$\$. In the radial and polar direction, we employed an outflow boundary condition, whereas the azimuthal boundary is periodic. There are $N_{r}=250$ cells in the radial, $N_{\theta}=100$ cells in the polar and $N_{\phi}=225$ cells in the azimuthal direction, resulting in $H/\Delta r=H/(r\Delta\theta)=H/(r\Delta\phi)\approx 1.3$ at $r=$5.2\text{\,}\mathrm{AU}$\$ (${\approx}3$ at $100\text{\,}\mathrm{AU}$). Here, $H$ is the gas pressure scale height and $\Delta r$, $\Delta\theta$ and $\Delta\phi$ are the resolution in the respective direction. We utilized an isothermal equation of state, $P=c_{s}^{2}\rho$, where $P$ is the gas pressure, $c_{s}$ is the isothermal sound speed, and $\rho$ is the gas density. A $1\text{\,}\mathrm{M_{\Sun}}$ star was placed at the center of the grid. Additionally, we add artificial viscosity to the simulation, corresponding to an $\alpha$ parameter (Shakura & Sunyaev 1973) of $\alpha=${10}^{-3}$\$. Due to the lack of self-gravity in our simulation, we deliberately omitted indirect terms rooted in the stellar acceleration caused by the gas in the potential.
@@ 2.2 Setup of the environment
We simulated the late accretion of gas onto Class II disks in two different scenarios. First, we considerd a controlled case where the environmental interactions are represented by a single, spherical cloudlet of gas encountering the disk on a hyperbolic orbit. Subsequently, we considerd the more realistic case of a disk with a systemic velocity embedded in a turbulent interstellar medium (ISM), utilizing a compressively turbulent initial condition with a given power spectrum and dispersion.
@@ 2.2.1 Spherical cloudlet
with $\nu$ the true anomaly. Furthermore, the initial velocity is given by
@@
The cloudlet was initialized with $v_{\infty}=$0.5\text{\,}\mathrm{km}\text{\,}{\mathrm{s}}^{-1}$\$, $d_{0}=$3000\text{\,}\mathrm{AU}$\$, a cloudlet radius of $R_{\mathrm{cloud}}=$500\text{\,}\mathrm{AU}$\$, and a cloudlet mass of $M_{\mathrm{cloud}}=$5\text{\times}{10}^{-3}\text{\,}\mathrm{M_{\Sun}}$=0.1M_{d}$, resulting in an initial density of $\rho_{\mathrm{cloud,0}}=$5.67\text{\times}{10}^{-18}\text{\,}\mathrm{g}\text{\,}{\mathrm{cm}}^{-3}$\$. We chose $b=0.5b_{\mathrm{crit}}$ for the impact parameter. The mass ratio was chosen such that the disk disruption due to the cloudlet is moderate. We selected the other parameter values with the goal to keep the radial domain of the grid and the physical time until the interaction small, and ensure $b<b_{\mathrm{crit}}$ across the entire physical extent of the cloudlet. The mass of the cloudlet is higher than suggested by observational data from giant molecular clouds (Falgarone et al. 2004; Klessen & Hennebelle 2010) for our choice of initial radius, but the lack of pressure support from a warm ISM leads to expansion of the cloudlet before encountering the disk. This initial condition is not intended to directly represent realistic physical conditions. Rather, we started with a more compact cloudlet to keep the inaccuracy incurred by neglecting the physical extent when initializing the orbital velocity small. When initially encountering the disk, the mean density of the cloudlet gas is ${\sim}${10}^{-19}\text{\,}\mathrm{g}\text{\,}{\mathrm{cm}}^{-3}$\$, which is comparable to the density of filamentary structures in the Taurus star forming region (Pineda et al. 2010; Palmeirim et al. 2013). This initial expansion does not affect the angular momentum of the cloudlet relative to the disk. Therefore, we do not expect it to substantially affect our results. The total runtime of the cloudlet simulation is $42.7\text{\,}\mathrm{kyr}$, after which any arising structures have dissipated.
@@ 2.2.2 Turbulent interstellar medium
where we chose $\hat{\@vec{v}}_{\mathrm{sys}}=(0,\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}})$ as the unit vector for the direction of systemic infall to consider one realization of the expected random alignment between the disk and systemic velocity. The velocity initial condition is therefore described by the two parameters $\sigma_{\mathrm{turb}}$ and $v_{\mathrm{sys}}$, with the two parameters $k_{\mathrm{min}}$ and $k_{\mathrm{max}}$ describing the limits of the turbulence power spectrum. We note that the turbulence was only introduced as an initial condition, and we did not apply turbulent forcing to drive the turbulence dynamically. As a result, the simulation is no longer self-consistent after the gas from the outer grid edge has reached the disk. At the highest systemic velocity we consider, $v_{\mathrm{sys}}=$1\text{\,}\mathrm{km}\text{\,}{\mathrm{s}}^{-1}$\$, this occurs after $237\text{\,}\mathrm{kyr}$, which is much larger than the employed simulation runtime of $30\text{\,}\mathrm{kyr}$. Nevertheless, the lack of turbulent forcing results in a decrease of the turbulent velocity dispersion over time. In terms of the turbulence Mach number $\mathcal{M}$, a turbulent velocity diffusion of $\sigma_{k}=\{1,0.5,0.25,0.1\}\ $\mathrm{km}\text{\,}{\mathrm{s}}^{-1}$\$ corresponds to $\mathcal{M}=\{8,4,3.2,2.9\}$ at the beginning of the simulation, which has decayed down to $\mathcal{M}=\{6,3,2.9,2.8\}$ in the end.
@@ 3 Results
In this work, we investigated the emergence of streamer-like structures in two different scenarios. First, we considered the case of a single gas cloudlet being accreted onto the disk (see Sect. 2.2.1). The primary advantage of this scenario is that it is a controlled environment with a well-defined encounter, however, we lack realism. Second, we considered the more realistic case of the movement of a star-disk system through the dense, turbulent ISM (see Sect. 2.2.2), which inherently contains more randomness.
@@ 3.1 Cloudlet capture
As a result of this difference, the cloudlet has already expanded considerable at the time when it encounters the disk. This can be seen in the second panel of Fig. 1, which shows the time evolution of the surface density of and around the disk, as well as the flow lines of the gas. Indeed, in the given field of view, the original spherical geometry is fully lost, and it can be seen that diffuse gas engulfs the disk fully during the encounter. As a direct consequence, over-shooting mass is accumulated on the side of the disk opposite to the infall direction. This morphology is reminiscent of an accretion tail that is the result of converging flow lines caused by gravitational focussing. A key difference to such an accretion tail lies in the angular momentum the accumulated material has with respect to the star. The hyperbolic orbit of the cloudlet has an impact parameter of $b=0.5b_{\mathrm{crit}}$, and angular momentum is not lost in the expansion process; instead, it is reflected by the asymmetry of the density and velocity at the time of the encounter.
The gas that is accumulated in the accretion tail remains bound to the star, because $b<b_{\mathrm{crit}}$. Due to the angular momentum it has with respect to the disk, it falls back onto it in an arced shape. This ”fall-back streamer” is supported in its shape by the flow of material that is still inflowing from the original direction, creating a shock at the edge of the arc. The resulting streamer morphology can be seen in the third panel of Fig. 1. Due to being caused by the fall-back of material that was accumulated in the accretion tail, we find that the direction of the streamer differs from the original inflow direction.
However, the resulting streamer is short-lived. Only ${\sim}$8.5\text{\,}\mathrm{kyr}$\$ later, the surface density of the fall-back streamer has decreased in contrast considerably, as shown in the forth panel of Fig. 1. Subsequently, the distribution of mass in the simulation can no longer be recognized as a streamer, although left-over material serves as evidence of the encounter until it is eventually accreted on timescales beyond the runtime of the simulation.
@@
Furthermore, we created synthetic observations of the CO J 2-1 transition. The resulting moment 1 and peak brightness temperature maps are shown in Fig. 3. The streamer can be recognized dynamically, adding a recognizable non-Keplerian component to the moment 1 map, similar to systems with observed streamers like HL Tau and DG Tau (Garufi et al. 2022). Closer to the disk, the moment 0 signal of the shock fall-back streamer exceeds $600\text{\,}\mathrm{mJy}\text{\,}{\mathrm{arcsec}}^{-2}\text{\,}\mathrm{km}\text{\,}{\mathrm{s}}^{-1}$, and the arc visible in scattered light is also clearly visible in the peak brightness temperature map. In the case of DG Tau, the channel map morphology was explained using a model of a cloudlet confined by a warm medium (Hanawa et al. 2024); here, we find that the accretion of a more extended gas structure, if it has angular momentum relative to the star, can lead to similar results.

### v2

# Emergence of streamers in simulations of late infall

@@
Emergence of streamers in simulations of late infall
@@
(December 23, 2025)
Growing observational evidence suggests that Class II protoplanetary disks may undergo substantial interactions with their environment in the form of late infall. This mass inflow predominantly manifests itself in the form of so-called streamers: filaments and arcs of gas connecting large-scale, extended gas structures to disk scales. Prevalent late infall has far-reaching consequences for planet formation theory, challenging the long-standing treatment of evolved disks in isolation. In this work, we investigate the emergence of late-infall streamers in different formation scenarios, their morphology and multiplicity, as well as their dependence on environmental conditions. We conducted this investigation by performing 3D hydrodynamical simulation using the grid-based code FARGO3D, which we post-process to obtain synthetic observations using the Monte Carlo radiative transfer code RADMC3D. We find that, while a late infall event in the form of a single encounter with a ”cloudlet” of gas can produce a streamer via an interplay between the fallback of bound material and shocks, such features dissipate quickly, on a timescale of ${\sim}$10\text{\,}\mathrm{kyr}$\$. Furthermore, we find that streamers can also form naturally in a turbulent, dense environment without the need for such encounters, which could act to reconcile short-lived streamers with ubiquitous detection of these structures. Here, we find multiple co-existing streamers for a disk velocity relative to the interstellar medium of $v_{\mathrm{sys}}=$0.5\text{\,}\mathrm{km}\text{\,}{\mathrm{s}}^{-1}$\$ and a turbulent velocity dispersion of $\sigma_{\mathrm{turb}}=$0.5\text{\,}\mathrm{km}\text{\,}{\mathrm{s}}^{-1}$\$, in which case an angular momentum flux of $\dot{J}\sim${10}^{38}\text{\,}\mathrm{g}\text{\,}{\mathrm{cm}}^{2}\text{\,}{\mathrm{s}}^{-2}$\$ arises. We find considerable dependence of the streamer morphology on the environment, which may act as a utility to constrain the physical conditions of the gas surrounding planet-forming disk, and therefore the conditions under which planets form.
@@ 1 Introduction
Planet formation theory has evolved greatly in recent years, with more complete models now connecting the large range of size scales of dust grain growth, the formation of planetary cores, migration, and gas accretion (for a review, see Drążkowska et al. 2023). Despite the steady advancement, protoplanetary disks, the birthplaces of planets, are still largely treated as being isolated from their environment. After their initial formation, the dust budget is treated as fixed, being converted into planetary cores through internal processes. Issues remain in this framework; in particular, recent observations indicate that the dust mass budget of fully formed protoplanetary disks is insufficient to form the giant planets we observe (Manara et al. 2018).
@@
Understanding the mechanisms that can create streamer-like structures and shape their morphology, as well as identifying their potential origin, is vital for our understanding of the dynamical and mass-budget evolution of protoplanetary disks, with potential far-reaching consequences for planet formation as a whole. In this work, we investigate two different types of interactions of fully formed protoplanetary disks with their environment using hydrodynamical simulations. First, we probed how a single, well-characterized interaction with a cloudlet of gas can create various streamer-like observational signatures (Section 3.1). We then considered a disk that is embedded in the turbulent interstellar medium to examine structures that arise in the absence of a distinct source of infall (Section 3.2). While signatures of single interactions are not long-lived, those arising from constant interactions with the ISM could persist over longer time scales and would therefore be much more abundantly observable.
@@ 2 Methods
We performed 3D hydrodynamical simulations on a spherical grid with logarithmic spacing in the radial direction, and uniform spacing in the polar and azimuthal direction. The computations were performed using the FARGO3D code (Benítez-Llambay & Masset 2016), capable of GPU multiprocessing while utilizing the advection algorithm by Masset (2000). The radial extent of the grid was from $r_{\mathrm{min}}=$10\text{\,}\mathrm{AU}$\$ to $r_{\mathrm{max}}=$5000\text{\,}\mathrm{AU}$\$, except the simulation presented in Section 3.1, where $r_{\mathrm{min}}=$5\text{\,}\mathrm{AU}$\$. We moved the outer edge further outward to $r_{\mathrm{max}}=$50\,000\text{\,}\mathrm{AU}$\$ for simulations with a turbulent initial condition, described in later sections. In the polar direction, the range was from $\theta_{\mathrm{min}}=$10\text{\,}\mathrm{\SIUnitSymbolDegree}$\$ to $\theta_{\mathrm{max}}=$170\text{\,}\mathrm{\SIUnitSymbolDegree}$\$. In the radial and polar direction, we employed an outflow boundary condition, whereas the azimuthal boundary is periodic. There are $N_{r}=250$ cells in the radial, $N_{\theta}=100$ cells in the polar and $N_{\phi}=225$ cells in the azimuthal direction, resulting in $H/\Delta r=H/(r\Delta\theta)=H/(r\Delta\phi)\approx 1.3$ at $r=$5.2\text{\,}\mathrm{AU}$\$ (${\approx}3$ at $100\text{\,}\mathrm{AU}$). Here, $H$ is the gas pressure scale height and $\Delta r$, $\Delta\theta$, and $\Delta\phi$ are the resolution in the respective direction. We utilized an isothermal equation of state, $P=c_{s}^{2}\rho$, where $P$ is the gas pressure, $c_{s}$ is the isothermal sound speed, and $\rho$ is the gas density. A $1\text{\,}\mathrm{M_{\Sun}}$ star was placed at the center of the grid. Additionally, we added artificial viscosity to the simulation, corresponding to an $\alpha$ parameter (Shakura & Sunyaev 1973) of $\alpha=${10}^{-3}$\$. Due to the lack of self-gravity in our simulation, we deliberately omitted indirect terms rooted in the stellar acceleration caused by the gas in the potential.
@@ 2.2 Setup of the environment
We simulated the late accretion of gas onto Class II disks in two different scenarios. First, we considered a controlled case where the environmental interactions are represented by a single, spherical cloudlet of gas encountering the disk on a hyperbolic orbit. Subsequently, we considered the more realistic case of a disk with a systemic velocity embedded in a turbulent interstellar medium (ISM), utilizing a compressively turbulent initial condition with a given power spectrum and dispersion.
@@ 2.2.1 Spherical cloudlet
with $\nu$ the true anomaly. The initial velocity is given by
@@
The cloudlet was initialized with $v_{\infty}=$0.5\text{\,}\mathrm{km}\text{\,}{\mathrm{s}}^{-1}$\$, $d_{0}=$3000\text{\,}\mathrm{AU}$\$, a radius of $R_{\mathrm{cloud}}=$500\text{\,}\mathrm{AU}$\$, and a mass of $M_{\mathrm{cloud}}=$5\text{\times}{10}^{-3}\text{\,}\mathrm{M_{\Sun}}$=0.1M_{d}$, resulting in an initial density of $\rho_{\mathrm{cloud,0}}=$5.67\text{\times}{10}^{-18}\text{\,}\mathrm{g}\text{\,}{\mathrm{cm}}^{-3}$\$. We chose $b=0.5b_{\mathrm{crit}}$ for the impact parameter. The mass ratio was chosen such that the disk disruption due to the cloudlet is moderate. We selected the other parameter values with the goal to keep the radial domain of the grid and the physical time until the interaction small, and ensure $b<b_{\mathrm{crit}}$ across the entire physical extent of the cloudlet. The mass of the cloudlet is higher than suggested by observational data from giant molecular clouds (Falgarone et al. 2004; Klessen & Hennebelle 2010) for our choice of initial radius, but the lack of pressure support from a warm ISM leads to expansion of the cloudlet before encountering the disk. This initial condition was not intended to directly represent realistic physical conditions. Rather, we started with a more compact cloudlet to keep the inaccuracy incurred by neglecting the physical extent when initializing the orbital velocity small. When initially encountering the disk, the mean density of the cloudlet gas is ${\sim}${10}^{-19}\text{\,}\mathrm{g}\text{\,}{\mathrm{cm}}^{-3}$\$, which is comparable to the density of filamentary structures in the Taurus star forming region (Pineda et al. 2010; Palmeirim et al. 2013). This initial expansion does not affect the angular momentum of the cloudlet relative to the disk. Therefore, we do not expect it to substantially affect our results. The total runtime of the cloudlet simulation is $42.7\text{\,}\mathrm{kyr}$, after which any arising structures have dissipated.
@@ 2.2.2 Turbulent interstellar medium
where we chose $\hat{\@vec{v}}_{\mathrm{sys}}=(0,\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}})$ as the unit vector for the direction of systemic infall to consider one realization of the expected random alignment between the disk and systemic velocity. The velocity initial condition is therefore described by the two parameters $\sigma_{\mathrm{turb}}$ and $v_{\mathrm{sys}}$, with the two parameters $k_{\mathrm{min}}$ and $k_{\mathrm{max}}$ describing the limits of the turbulence power spectrum. We note that the turbulence was only introduced as an initial condition, and we did not apply turbulent forcing to drive the turbulence dynamically. As a result, the simulation is no longer self-consistent after the gas from the outer grid edge has reached the disk. At the highest systemic velocity we considered, $v_{\mathrm{sys}}=$1\text{\,}\mathrm{km}\text{\,}{\mathrm{s}}^{-1}$\$, this occurs after $237\text{\,}\mathrm{kyr}$, which is much larger than the employed simulation runtime of $30\text{\,}\mathrm{kyr}$. Nevertheless, the lack of turbulent forcing results in a decrease of the turbulent velocity dispersion over time. In terms of the turbulence Mach number $\mathcal{M}$, a turbulent velocity diffusion of $\sigma_{k}=\{1,0.5,0.25,0.1\}\ $\mathrm{km}\text{\,}{\mathrm{s}}^{-1}$\$ corresponds to $\mathcal{M}=\{8,4,3.2,2.9\}$ at the beginning of the simulation, which has decayed down to $\mathcal{M}=\{6,3,2.9,2.8\}$ in the end.
@@ 3 Results
In this work, we investigated the emergence of streamer-like structures in two different scenarios. First, we considered the case of a single gas cloudlet being accreted onto the disk (see Sect. 2.2.1). The primary advantage of this scenario is that it is a controlled environment with a well-defined encounter, however, we lack realism. Second, we considered the more realistic case of a star-disk system moving through the dense, turbulent ISM (see Sect. 2.2.2), which inherently contains more randomness.
@@ 3.1 Cloudlet capture
As a result of this difference, the cloudlet has already expanded considerably at the time when it encounters the disk. This is apparent in the second panel of Fig. 1, showing the time evolution of the surface density in and around the disk, as well as the flow lines of the gas. Indeed, in the given field of view, the original spherical geometry is fully lost, and it can be seen that diffuse gas engulfs the disk fully during the encounter. As a direct consequence, over-shooting mass is accumulated on the side of the disk opposite to the infall direction. This morphology is reminiscent of an accretion tail that is the result of converging flow lines caused by gravitational focussing. A key difference to such an accretion tail lies in the angular momentum the accumulated material has with respect to the star. The hyperbolic orbit of the cloudlet has an impact parameter of $b=0.5b_{\mathrm{crit}}$, and angular momentum is not lost in the expansion process; instead, it is reflected by the asymmetry of the density and velocity at the time of the encounter.
The gas that is accumulated in the accretion tail remains bound to the star, because $b<b_{\mathrm{crit}}$. Due to the angular momentum it has with respect to the disk, it falls back onto it in an arced shape. This ”fallback streamer” is supported in its shape by the flow of material that is still inflowing from the original direction, creating a shock at the edge of the arc. The resulting streamer morphology can be seen in the third panel of Fig. 1. Due to being caused by the fallback of material that was accumulated in the accretion tail, we find that the direction of the streamer differs from the original inflow direction.
However, the resulting streamer is short-lived. Only ${\sim}$8.5\text{\,}\mathrm{kyr}$\$ later, the surface density of the fallback streamer has decreased in contrast considerably, as shown in the forth panel of Fig. 1. Subsequently, the distribution of mass in the simulation can no longer be recognized as a streamer, although left-over material serves as evidence of the encounter until it is eventually accreted on timescales beyond the runtime of the simulation.
@@
Furthermore, we created synthetic observations of the CO J 2-1 transition. The resulting moment 1 and peak brightness temperature maps are shown in Fig. 3. The streamer can be recognized dynamically, adding a recognizable non-Keplerian component to the moment 1 map, similar to systems with observed streamers like HL Tau and DG Tau (Garufi et al. 2022). Closer to the disk, the moment 0 signal of the shock fallback streamer exceeds $600\text{\,}\mathrm{mJy}\text{\,}{\mathrm{arcsec}}^{-2}\text{\,}\mathrm{km}\text{\,}{\mathrm{s}}^{-1}$, and the arc visible in scattered light is also clearly visible in the peak brightness temperature map. In the case of DG Tau, the channel map morphology was explained using a model of a cloudlet confined by a warm medium (Hanawa et al. 2024); here, we find that the accretion of a more extended gas structure, if it has angular momentum relative to the star, can lead to similar results.